Calculate the first cosmic velocity for the planet Jupiter. The radius of Jupiter is R = 7 × 10⁴ km, while the acceleration of free fall near the surface of the planet is g = 26m / s².


1
Expert's answer
2021-01-20T14:58:19-0500

We can find the first cosmic velocity from the formula:


"v=\sqrt{\dfrac{GM}{R}},""v=\sqrt{\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot1.89\cdot10^{27}\ kg}{7\cdot10^7\ m}}=42.4\ \dfrac{km}{s}."

Answer:

"v=42.4\ \dfrac{km}{s}."